OpenStudy (anonymous):

A carpet has a rectangular measure of 10 feet for the length and w feet for the width. If the perimeter is the sum of all four sides of the field, what is an expression for the perimeter of the field? Simplify.

OpenStudy (godlovesme):

p=2(l+w)

OpenStudy (godlovesme):

make sense @lexi98

OpenStudy (kamibug):

And we could plug in 10 for l since that's given. p = 2(10+w) We're told to simplify as well. That would be p = 20 + 2w \(\Huge\cal\color{royalblue}{Welcome~to~Openstudy!~:)}\)

OpenStudy (anonymous):

Thank you @KamiBug

OpenStudy (kamibug):

No problem. :)